Juarez, Inc. uses a job-order costing system for its products, which pass from the Machining Department, to the Assembly Department, to finished-goods inventory. The Machining Department is heavily automated; in contrast, the Assembly Department performs a number of manual-assembly activities. The company applies manufacturing overhead using machine hours in the Machining Department and direct-labor cost in the Assembly Department. The following information relates to the year just ended:

Machining Department Assembly Department
  Budgeted manufacturing overhead $4,000,000 $3,136,000
  Actual manufacturing overhead 4,270,000 3,030,000
  Budgeted direct-labor cost (based on practical capacity) 1,500,000 5,600,000
  Actual direct-labor cost 1,450,000 5,780,000
  Budgeted machine hours (based on practical capacity) 400,000 100,000
  Actual machine hours 425,000 110,000
The data that follow pertain to job no. DC66, the only job in production at year-end.
Machining Department Assembly Department
  Direct material $24,500 $ 6,700
  Direct labor $27,800 $58,700
  Machine hours 360 150
Selling and administrative expense amounted to $2,500,000.
5. If Juarez disposes of under- or overapplied overhead as an adjustment to Cost of Goods Sold, would the company’s Cost of Goods Sold account increase or decrease?
6.

How much overhead would have been charged to the company’s Work-in-Process account during the year?

7. Comment on the appropriateness of the company’s cost drivers (i.e., the use of machine hours in Machining and direct-labor cost in Assembly).
- The company’s cost drivers are not appropriate. OR
- The company’s cost drivers are appropriate.
Answer #1
Machining Department Assembly Department
Budgeted manufacturing overhead 4000000 3136000
Divide by Cost driver 400000 5600000
Predetermined overhead rate 10 56%
Overhead applied:
Machining Department 4250000 =425000*10
Assembly Department 3236800 =5780000*56%
Overhead applied 7486800
Actual manufacturing overhead 7300000 =4270000+3030000
5
Cost of Goods Sold account will decrease as Overhead is over-applied
6
Overhead charged to Work-in-Process account = $7486800
7
The company’s cost drivers are appropriate.